a certain radioactive substance has a half-life of 40 minutes. How much of an original 100 grams substance will remain after 2 hours?

Grade:12th Pass

2 Answers

VEERU YERRAYYA
11 years ago

half life = 40minutes =2/3 hours

t(1/2)=(ln2 /k)

then k=(ln2/t(1/2)) = ln2/(2/3) = 3/2*(ln2)

origianl substance y0= 1000 grams

y=y0e-kt

y= 1000*(e-(3/2(ln2))*2 ) = 1000/8 = 125 grams

bhaveen kumar
11 years ago

m= m0/2t/T

m=100 gms..

t=2 hours=2*60=120 min..

T=40 min

m=100/2120/40

m=100/23

m=100/8

m=12.5 gms.....